# Deep Compression Autoencoder for Efficient High-Resolution Diffusion Models

### v1

@@ 2 Related Work
Training and evaluating diffusion models directly in high-resolution pixel space results in prohibitive computational costs. To address this issue, Rombach et al. (2022) proposes latent diffusion models that operate in a compressed latent space produced by pretrained autoencoders. The proposed autoencoder with 8×8$\times$8 × spatial compression ratio and $4$ latent channels has been widely adopted in subsequent works (Peebles & Xie, 2023; Bao et al., 2023). Since then, follow-up works mainly focus on enhancing the reconstruction accuracy of the f8 autoencoder by increasing the number of latent channels (Esser et al., 2024; Dai et al., 2023; Labs, 2024). In contrast, our work focuses on an orthogonal direction, increasing the spatial compression ratio of the autoencoders (e.g., f64). To the best of our knowledge, our work is the first study in this critical but underexplored direction.
@@
Diffusion models have been widely used for image generation and showed impressive results (Labs, 2024; Esser et al., 2024). However, diffusion models are computationally intensive, motivating many works to accelerate diffusion models. One representative strategy is reducing the number of inference sampling steps by training-free few-step samplers (Song et al., 2021; Lu et al., 2022a; b; Zheng et al., 2023; Zhang & Chen, 2023; Zhang et al., 2023; Zhao et al., 2024b; Shih et al., 2024; Tang et al., 2024) or distilling-based methods (Meng et al., 2023; Salimans & Ho, 2022; Yin et al., 2024b; a; Song et al., 2023; Luo et al., 2023; Liu et al., 2023). Another representative strategy is model compression by leveraging sparsity (Li et al., 2022; Ma et al., 2024) or quantization (He et al., 2024; Fang et al., 2024; Li et al., 2023; Zhao et al., 2024a). Designing efficient diffusion model architectures (Li et al., 2024c; Liu et al., 2024; Cai et al., 2024) or inference systems (Li et al., 2024b; Wang et al., 2024) is also an effective approach for boosting efficiency. In addition, improving the data quality (Chen et al., 2024b; a) can boost the training efficiency of diffusion models.
@@ 3.1 Motivation
Additionally, we increase the latent channel number to maintain the same total latent size across different settings. We can then convert the latent to a higher spatial compression ratio one by applying a space-to-channel operation: $H\times W\times C\rightarrow\frac{H}{p}\times\frac{W}{p}\times p^{2}C$.
@@ 3.3 Application to Latent Diffusion Models
We conduct ablation study experiments and summarize the results in Table 1. We can see that directly reaching the target spatial compression ratio with the autoencoder gives the best results among all settings. In addition, we also find that shifting the spatial compression ratio from the diffusion model to the autoencoder consistently leads to better FID. We conjecture it is because the latent diffusion models need to simultaneously learn denoising and token compression when using a patch size $>1$. With the autoencoder taking over the whole token compression task, the diffusion model can fully focus on the denoising task and thus achieves better results.
@@ 4.1 Setups
For image generation experiments, we apply autoencoders to diffusion transformer models including DiT (Peebles & Xie, 2023) and UViT (Bao et al., 2023). We follow the same training settings as the original papers. We consider three settings with different resolutions, including ImageNet (Deng et al., 2009) for $512\times 512$ generation, FFHQ (Karras et al., 2019) and MJHQ (Li et al., 2024a) for $1024\times 1024$ generation, and MapillaryVistas (Neuhold et al., 2017) for $2048\times 2048$ generation.
@@
Table 3: Class-Conditional Image Generation Results on ImageNet 512$\times$512. represents the model is trained for 4$\times$ training iterations (i.e., 500K $\rightarrow$ 2, [^0] K iterations).
2 84 833 14.1 41.2 16.98 48.05 38.19
SD-VAE-f8 [30] 4 470 5566 2.5 10.7 23.81 60.94 51.29
DC-AE-f32 1 475 5575 2.5 10.7 13.65 34.35 27.20
DiT-S [29] DC-AE-f64 1 2085 25259 1.0 3.1 26.88 61.30 53.38
MapillaryVistas 2048$\times$2048 (Unconditional)
SD-VAE-f8 [30] 4 84 810 14.3 41.4 69.50
DiT-S [29] DC-AE-f64 1 459 5435 2.6 11.0 59.55
Table 4: 1024$\times$1024 and 2048$\times$2048 Image Generation Results.
@@ 4.3 Latent Diffusion Models
We compare DC-AE with the widely used SD-VAE-f8 autoencoder (Rombach et al., 2022) on various diffusion transformer models. For DC-AE, we always use a patch size of 1 (denoted as p1). For SD-VAE-f8, we follow the common setting and use a patch size of 2 or 4 (denoted as p2, p4). The results are summarized in Table 3, Table 4, and Table 5.
@@
As shown in Table 3, DC-AE-f32p1 consistently delivers better FID than SD-VAE-f8p2 on all diffusion transformer models. In addition, it has 4$\times$ fewer tokens than SD-VAE-f8p2, leading to 4.5$\times$ higher H100 training throughput and 4.8$\times$ higher H100 inference throughput for DiT-XL. We also observe that larger diffusion transformer models seem to benefit more from our DC-AE. For example, DC-AE-f64p1 has a worse FID than SD-VAE-f8p2 on UViT-S but a better FID on UViT-H. We conjecture it is because DC-AE-f64 has a larger latent channel number than SD-VAE-f8, thus needing more model capacity (Esser et al., 2024).
1024$\times$1024 and 2048$\times$2048 Image Generation.

### v4

@@ 2 Related Work
Training and evaluating diffusion models directly in high-resolution pixel space results in prohibitive computational costs. To address this issue, Rombach et al. (2022) proposes latent diffusion models that operate in a compressed latent space produced by pretrained autoencoders. The proposed autoencoder with 8×8$\times$8 × spatial compression ratio and $4$ latent channels has been widely adopted in subsequent works (Peebles & Xie, 2023; Bao et al., 2023). Since then, follow-up works mainly focus on enhancing the reconstruction accuracy of the f8 autoencoder by increasing the number of latent channels (Esser et al., 2024; Dai et al., 2023; Labs, 2024). Additionally, to improve the reconstruction quality, Zhu et al. (2023) leverages a heavier decoder and incorporates task-specific priors. In contrast to prior works, our work focuses on an orthogonal direction, increasing the spatial compression ratio of the autoencoders (e.g., f64). To the best of our knowledge, our work is the first study in this critical but underexplored direction.
@@
Diffusion models have been widely used for image generation and showed impressive results (Labs, 2024; Esser et al., 2024). However, diffusion models are computationally intensive, motivating many works to accelerate diffusion models. One representative strategy is reducing the number of inference sampling steps by training-free few-step samplers (Song et al., 2021; Lu et al., 2022a; b; Zheng et al., 2023; Zhang & Chen, 2023; Zhang et al., 2023; Zhao et al., 2024b; Shih et al., 2024; Tang et al., 2024) or distilling-based methods (Meng et al., 2023; Salimans & Ho, 2022; Yin et al., 2024b; a; Song et al., 2023; Luo et al., 2023; Liu et al., 2023). Another representative strategy is model compression by leveraging sparsity (Li et al., 2022; Ma et al., 2024b) or quantization (He et al., 2024; Fang et al., 2024; Li et al., 2023; Zhao et al., 2024a). Designing efficient diffusion model architectures (Li et al., 2024d; Liu et al., 2024; Cai et al., 2024) or inference systems (Li et al., 2024b; Wang et al., 2024) is also an effective approach for boosting efficiency. In addition, improving the data quality (Chen et al., 2024b; a) can boost the training efficiency of diffusion models.
@@ 3.1 Motivation
Additionally, we increase the latent channel number to maintain the same total latent size across different settings. We can then convert the latent to a higher spatial compression ratio one by applying a space-to-channel operation (Shi et al., 2016): $H\times W\times C\rightarrow\frac{H}{p}\times\frac{W}{p}\times p^{2}C$.
@@ 3.3 Application to Latent Diffusion Models
We conduct ablation study experiments and summarize the results in Table 1. We can see that directly reaching the target spatial compression ratio with the autoencoder gives the best results among all settings. In addition, we also find that shifting the spatial compression ratio from the diffusion model to the autoencoder consistently leads to better FID.
@@ 4.1 Setups
For image generation experiments, we apply autoencoders to diffusion transformer models including DiT (Peebles & Xie, 2023) and UViT (Bao et al., 2023). We follow the same training settings as the original papers. Additionally, we build USiT by combining UViT (Bao et al., 2023) with the SiT sampler (Ma et al., 2024a). We consider three settings with different resolutions, including ImageNet (Deng et al., 2009) for $512\times 512$ generation, FFHQ (Karras et al., 2019) and MJHQ (Li et al., 2024a) for $1024\times 1024$ generation, and MapillaryVistas (Neuhold et al., 2017) for $2048\times 2048$ generation.
@@
Table 3: Class-Conditional Image Generation Results on ImageNet 512$\times$512. represents the model is trained for 4$\times$ training iterations (i.e., 500K $\rightarrow$ 2, [^0] K iterations). represents the model is trained with 4$\times$ batch size (i.e., 256 $\rightarrow$ 1024).
@@ 4.2 Image Compression and Reconstruction
Figure 9: Model Scaling Results on ImageNet 512$\times$512 with UViT. DC-AE-f64 benefits more from scaling up than SD-VAE-f8.
@@ 4.3 Latent Diffusion Models
We compare DC-AE with the widely used SD-VAE-f8 autoencoder (Rombach et al., 2022) on various diffusion transformer models. For DC-AE, we always use a patch size of 1 (denoted as p1). For SD-VAE-f8, we follow the common setting and use a patch size of 2 or 4 (denoted as p2, p4). The results are summarized in Table 3, Table 4, and Figure 9.
@@
As shown in Table 3, DC-AE-f32p1 consistently delivers better FID than SD-VAE-f8p2 on all diffusion transformer models. In addition, it has 4$\times$ fewer tokens than SD-VAE-f8p2, leading to 4.5$\times$ higher H100 training throughput and 4.8$\times$ higher H100 inference throughput for DiT-XL. We also observe that larger diffusion transformer models seem to benefit more from our DC-AE (Figure 9). For example, DC-AE-f64p1 has a worse FID than SD-VAE-f8p2 on UViT-S but a better FID on UViT-2B. We conjecture it is because DC-AE-f64 has a larger latent channel number than SD-VAE-f8, thus needing more model capacity (Esser et al., 2024).
Applying DC-AE to USiT models, we achieve highly competitive results compared with prior leading image generative models. For example, DC-AE-f32+USiT-2B achieves 1.72 FID on ImageNet 512$\times$512, outperforming the SOTA diffusion model EDM2-XXL and SOTA auto-regressive image generative models (MAGVIT-v2 and MAR-L).
@@ Appendix A Additional Samples
Appendix B Diffusion Model Architecture Details
In addition to existing UViT models, we scaled the model up to 1.6B parameters, with a depth of 28, a hidden dimension of 2048, and 32 heads. We denote this model as UViT-2B.
Appendix C High-Resolution Image Generation Results
SD3-VAE-f8 [9] 2 83 814 14.2 41.4 46.28 109.43 103.02
Flux-VAE-f8 [18] 2 83 814 14.2 41.4 59.15 143.16 139.06
SDXL-VAE-f8 [36] 2 84 833 14.1 41.2 16.82 49.00 39.21
Asym-VAE-f8 [54] 2 84 833 14.1 41.2 17.12 48.25 38.36
2 84 833 14.1 41.2 16.98 48.05 38.19
SD-VAE-f8 [37] 4 470 5566 2.5 10.7 23.81 60.94 51.29
DC-AE-f32 1 475 5575 2.5 10.7 13.65 34.35 27.20
  DC-AE-f32 1 475 5575 2.5 10.7 11.39 28.36 21.89
DiT-S [35] DC-AE-f64 1 2085 25259 1.0 3.1 26.88 61.30 53.38
MapillaryVistas 2048$\times$2048 (Unconditional)
SD-VAE-f8 [37] 4 84 810 14.3 41.4 69.50
DiT-S [35] DC-AE-f64 1 459 5435 2.6 11.0 59.55
Table 5: 1024$\times$1024 and 2048$\times$2048 Image Generation Results. represents the model is trained with 4$\times$ batch size (i.e., 256 $\rightarrow$ 1024).
Appendix D Image Generation Results with Other Evaluation Metrics
Diffusion Patch Inference FID $\downarrow$ Inception Score $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ CMMD $\downarrow$
Model Autoencoder Size Throughput w/o CFG w/ CFG w/o CFG w/ CFG w/o CFG w/ CFG w/o CFG w/ CFG w/o CFG w/ CFG
SD3-VAE-f8 [9] 2 2984 164.34 143.82 6.07 7.53 0.06 0.09 0.31 0.39 3.13 2.94
Flux-VAE-f8 [18] 2 2984 106.07 84.73 13.39 17.71 0.28 0.37 0.39 0.42 1.90 1.67
SDXL-VAE-f8 [36] 2 2991 51.03 26.38 27.58 56.72 0.57 0.74 0.58 0.50 1.35 1.05
Asym-VAE-f8 [54] 2 2991 52.68 25.14 30.22 65.27 0.58 0.74 0.62 0.51 1.09 0.80
SD-VAE-f8 [37] 2 2991 51.96 24.57 30.37 65.73 0.57 0.74 0.64 0.52 1.23 0.91
SD-VAE-f16 [37] 2 12881 76.86 44.22 21.38 43.35 0.43 0.62 0.60 0.55 1.83 1.46
SD-VAE-f32 [37] 1 12883 70.23 38.63 23.07 47.72 0.46 0.64 0.58 0.56 1.71 1.36
DC-AE-f32 1 12850 46.12 18.08 34.82 84.73 0.59 0.76 0.66 0.56 1.00 0.70
DC-AE-f64 1 53774 67.30 35.96 24.55 52.86 0.44 0.64 0.60 0.56 1.44 1.14
UViT-S [1]   DC-AE-f64 1 53774 61.84 30.63 27.28 61.76 0.47 0.67 0.63 0.56 1.35 1.04
Flux-VAE-f8 [18] 2 416 27.35 8.72 53.09 130.20 0.68 0.83 0.61 0.48 0.54 0.30
Asym-VAE-f8 [54] 2 424 11.39 2.97 108.70 241.10 0.75 0.83 0.65 0.53 0.37 0.20
SD-VAE-f8 [37] 2 424 12.03 3.04 105.25 240.82 0.75 0.84 0.64 0.54 0.43 0.25
DC-AE-f32 1 2016 9.56 2.84 117.49 226.98 0.75 0.82 0.64 0.55 0.34 0.22
DiT-XL [35]   DC-AE-f32 1 2016 6.88 2.41 141.07 263.56 0.76 0.82 0.63 0.56 0.29 0.18
Flux-VAE-f8 [18] 2 349 30.91 12.63 56.72 127.93 0.64 0.76 0.59 0.49 0.50 0.31
Asym-VAE-f8 [54] 2 351 11.36 3.51 124.24 249.21 0.75 0.82 0.61 0.53 0.32 0.20
SD-VAE-f8 [37] 2 351 11.04 3.55 125.08 250.66 0.75 0.82 0.61 0.53 0.39 0.26
DC-AE-f32 1 1622 9.83 2.53 121.91 255.07 0.76 0.83 0.65 0.54 0.34 0.20
DC-AE-f64 1 6706 13.96 3.01 99.20 229.16 0.73 0.83 0.64 0.53 0.50 0.31
UViT-H [1] DC-AE-f64 1 6706 12.26 2.66 109.20 239.82 0.73 0.82 0.67 0.57 0.43 0.27
Flux-VAE-f8 [18] 2 155 25.03 10.12 74.04 161.29 0.67 0.78 0.58 0.51 0.38 0.24
Asym-VAE-f8 [54] 2 157 9.87 3.62 131.95 258.63 0.76 0.83 0.59 0.52 0.30 0.19
SD-VAE-f8 [37] 2 157 9.73 3.57 132.86 260.50 0.76 0.83 0.59 0.52 0.37 0.24
DC-AE-f32 1 665 8.13 2.30 135.44 272.73 0.76 0.82 0.66 0.56 0.30 0.17
DC-AE-f64 1 2733 7.78 2.47 138.11 280.49 0.77 0.84 0.63 0.54 0.35 0.22
UViT-2B [1] DC-AE-f64 1 2733 6.50 2.25 152.35 293.45 0.77 0.83 0.65 0.56 0.31 0.19
USiT-H DC-AE-f32 1 1622 3.80 1.89 174.58 252.35 0.78 0.82 0.64 0.60 0.24 0.18
USiT-2B DC-AE-f32 1 665 2.90 1.72 187.68 248.10 0.79 0.82 0.63 0.61 0.21 0.17
Table 6: Class-Conditional Image Generation Results on ImageNet 512$\times$512 with More Evaluation Metrics. represents the model is trained for 4$\times$ training iterations (i.e., 500K $\rightarrow$ 2, [^0] K iterations). represents the model is trained with 4$\times$ batch size (i.e., 256 $\rightarrow$ 1024).
Table 6 presents a comprehensive evaluation of different diffusion models and autoencoders on ImageNet 512$\times$512. The evaluation metrics include FID (Martin et al., 2017), inception score (IS) (Salimans et al., 2016), precision, recall (Kynkäänniemi et al., 2019), and CMMD (Jayasumana et al., 2024). Our DC-AE consistently delivers significant efficiency improvements while maintaining the generation performance under different evaluation metrics.